# DiffusionTrend: A Minimalist Approach to Virtual Fashion Try-On

### v1

@@ 3.3 Virtual Try-On Image Reconstruction
TryOnDiffusion Zhu et al. (2023) - - 13.447 6.964 - - - -
DCI-VTON Gou et al. (2023) 0.0530 0.8920 9.130 0.870 0.0443 - 11.800 -
LaDI-VTON Morelli et al. (2023) 0.0910 0.8760 9.410 0.160 0.0640 0.9060 6.480 0.220
WarpDiffusion Li et al. (2023a) 0.0880 0.9850 8.610 - 0.0890 0.9010 9.187 -
OOTDiffusion Xu et al. (2024) 0.0710 0.8780 8.810 0.820 0.0450 0.9270 4.200 0.370
StableVITON Kim et al. (2023) 0.0732 0.8880 8.233 0.490 0.0388 0.9370 9.940 0.120
IDM-VTON Choi et al. (2024) 0.1020 0.8700 6.290 - 0.0620 0.9200 8.640 -
Wear-Any-Way Chen et al. (2024) 0.0780 0.8770 8.155 0.780 0.0409 0.9340 11.720 0.330
DiffusionTrend (Ours) 0.0919 0.8575 10.864 0.540 0.0721 0.9170 9.856 0.430
4 Experimentation
4.1 Experimental Setup
Dataset. We conduct extensive experiments on two high-resolution datasets from the VITON benchmark: VITON-HD Choi et al. (2021) and DressCode Morelli et al. (2022). The VITON-HD dataset includes 13,679 pairs of images, each featuring a front-view upper-body shot of women alongside corresponding in-store garments, split into 11,647 training pairs and 2,032 testing pairs. The DressCode dataset is larger, with 48,392 training pairs and 5,400 testing pairs, featuring front-view full-body images of individuals with corresponding in-store garments, categorized into upper-body, lower-body, and dresses.
@@ 4.1 Experimental Setup
Evaluation Metrics. We assess both paired and unpaired scenarios. In the paired scenario, the target human image and its corresponding garment image are used for reconstruction. In the unpaired scenario, different garment images are used for the virtual try-on experience. To evaluate the quality of images generated in the paired scenario, we use LPIPS Zhang et al. (2018) and SSIM Wang et al. (2004) metrics to measure resemblance to the original image. In the unpaired scenario, we use FID Heusel et al. (2017) and KID Bińkowski et al. (2018) to gauge the realism and fidelity of the synthesized images.
Implementation Details. We use the Adam optimizer Kingma and Ba (2014) for the apparel localization network with a learning rate of 1e-4, halving it every 10 epochs. The network is trained for 35 epochs on two RTX 3090 GPUs. During inference, the number of clusters $K$ is set to 5. We set the apparel-related prompt to “clothes” and apply our method to Stable Diffusion XL (SDXL) Podell et al. (2023). For the inversion phase, we use an empty prompt, and for model image generation, the prompt is “model wearing clothes.” We conduct DDIM sampling Song et al. (2020) with 50 steps and set the classifier-free guidance to 7.5. Garment latent infusion occurs at time step $t_{1}=40$, and background restoration at $t_{2}=15$. The entire generation process is carried out on a single RTX 3090 GPU.
Figure 7: Qualitative comparisons on VITON-HD Choi et al. (2021)(1st $\sim$ 2nd row) and DressCode Morelli et al. (2022) (3rd $\sim$ 8th row).
4.2 Experimental Results
Quantitative Results. Table 2 shows the quantitative comparisons between DiffusionTrend and other methods on VITON-HD Choi et al. (2021) and DressCode Morelli et al. (2022) test datasets. The results on DressCode of DCI-VTON Gou et al. (2023) and StableVITON Kim et al. (2023) contain our implementation because no usable codes are given. Despite relying solely on pre-trained models for tasks like garment latent infusion and background restoration, our performance lags behind SOTA models. Past methods incur high training costs to ensure generated try-on images remain realistic and natural under various complex poses, as confirmed by evaluation metrics. In contrast, DiffusionTrend aims to provide a resource-efficient, user-friendly tool for quickly confirming purchase intentions under simple try-on poses. Without training the diffusion model on large datasets, our basic operations fall short in handling complex poses, leading to suboptimal performance. As the first visual try-on model without training on diffusion models, our approach differs from traditional methods in motivation and technical approaches, making traditional try-on dataset evaluations insufficient for comprehensively measuring our method’s effectiveness. Thus, quantitative experiments serve only as a reference, while qualitative experiments will further demonstrate our superiority.
@@ 4.2 Experimental Results
Qualitative Results. Fig. 7 provides the qualitative comparison of DiffusionTrend with the state-of-the-art baselines on the VITON-HD Choi et al. (2021) and DressCode Morelli et al. (2022) datasets. The results indicate that our DiffusionTrend performs as well as baseline methods under simple poses.
@@
Finally, our method is not limited by the training dataset and works well on different types of images.
@@
Our results, as presented in Table 3, demonstrate that our compact CNN achieves high accuracy in mask extraction tasks, with IoU scores reaching 0.7 and Dice scores exceeding 0.8. While its accuracy may lag behind SOTA segmentation models, it is more than sufficient for handling virtual try-on tasks. These results validate the effectiveness and practicality of our approach.
Table 3: Quantitative metrics on virtual try-on datasets for validating mask prediction accuracy.
@@
4.3 Ablation Study
In this section, we examine the optimality of our method’s components and experimental settings, including background restoration and the timestep for latent infusion or replacement in Sec. 3.3. Qualitative and quantitative results are presented in Fig. 8 and Table 4, respectively.
Table 4: Quantitative ablations for $t_{1}$ in Garment Latent Infusion and $t_{2}$ in Background Restoration. Bold numbers indicate the best performance.
@@ 4.3 Ablation Study
      50       15       0.0804       0.9018       10.1231       0.44
      45       0.0763       0.9036       9.9633       0.44
      40 (Ours)       0.0731       0.9049       9.9322       0.44
      35       0.0722       0.9053       10.0390       0.46
      30       0.0709       0.9060       10.2479       0.47
      25       0.0701       0.9064       10.4147       0.48
      40       25       0.0765       0.9024       9.4499       0.41
      20       0.0748       0.9037       9.6527       0.43
      15 (Ours)       0.0731       0.9049       9.9322       0.44
      10       0.0735       0.9050       10.2813       0.46
      5       0.0736       0.9050       10.9275       0.51
      0       0.1697       0.8404       16.1841       0.76
Figure 9: Color oversaturation issue can be alleviated by reducing the sampling steps.
@@
$t_{1}$ for Garment Latent Infusion. The upper part of Table 4 shows that incorporating clothing information too early tends to lower the LPIPS, SSIM, and FID scores. Specifically, as observed in the top row of Fig. 8, prematurely incorporating garment information results in a deficiency of integration between the model and the garment, effectively reconstructing their respective latent representations in disparate areas without any interaction. This issue is clearly visible, as there are pronounced boundaries between the clothing and the human figure, creating an impression of disjunction rather than a cohesive unity. On the other hand, introducing the information at a later stage impedes the development and enhancement of the garment’s finer details. This is observable in the figure, where the cartoon patterns are prone to becoming indistinct, thereby compromising the overall quality of the image.
@@
$t_{2}$ for Background Restoration. Although the quantitative results in Table 4 show that the FID and KID scores are higher at $t_{2}=25$ and $20$, it can be observed from the second row of Fig. 8 that, performing background restoration too early can negatively impact the generation of garment details. For instance, at $t_{2}=25$, the waist’s metallic embellishments are lost during the reconstruction process; at $t_{2}=20$, the situation improves slightly, but the metallic dots on either side are still missing. In contrast, performing it too late creates a distinct boundary between the background and the foreground, resulting in unnatural outcomes, and leads to the over-rendering of details. In the figure at $t_{2}=15$, there are small red artifacts visible around the metallic dots.
5 Limitations and Future Work
@@
Subtle Variations in Visual Fidelity. The refinement in clothing generation is not always optimal, occasionally leading to minor color variations and the presence of unnatural textures and patterns. After extensive experimentation, we suggest mitigating it by appropriately reducing the sampling steps.
As demonstrated in Fig. 9, when the sampling steps are set to 46, the generated clothing colors more closely match the original garment image. This adjustment significantly improves the visual fidelity of the results. However, we do not recommend reducing the sampling steps too much, as we also discover that excessive reduction introduces new issues. For instance, when the steps are reduced to 44, we observe distortion and inconsistency in the details of both the person and the background, which degrades the overall visual quality.
@@
Additionally, some studies Chung et al. (2024) suggest that the color of generated images may be influenced by the classifier-free guidance scale. However, our attempts do not yield any significant improvements. Nevertheless, we will continue to explore other methods to further enhance the robustness of DiffusionTrend.
Reliance on Inversion Quality. Our proposed DiffusionTrend’s performance is indeed influenced by the quality of DDIM inversion results. However, we find that the evolution of more powerful diffusion models is likely to yield better results even with the same inversion method, As demonstrated in Fig. 10. Moreover, ongoing advancements in inversion methods, such as NULL-Text Inversion Mokady et al. (2023), ReNoise Garibi et al. (2024), Fixed-Point Iteration Pan et al. (2023), and EasyInv Zhang et al. (2024), provide promising directions for further improvements.
We believe that the emergence of these improved diffusion models and inversion techniques will significantly expand the application space of our training-free try-on paradigm.
@@
Inference Overhead and Optimization Prospects. DiffusionTrend requires additional time during inference due to the DDIM inversion process, which involves a similar number of steps as DDIM sampling. This time overhead mainly occurs during the inference process and is a potential drawback. However, it can be mitigated through quantization (e.g., Q-Diffusion Li et al. (2023b), NDTC Shang et al. (2023)) and pruning (e.g., LD-Pruner Castells et al. (2024)). While this paper focuses on reducing training overhead, combining these techniques to optimize inference speed will be an important direction for future research.
@@
6 Conclusion
In this paper, we have introduced DiffusionTrend, a novel try-on methodology that forgoes the need for training diffusion models, thereby offering straightforward, conventional pose virtual try-on services with minimal computational demands. Capitalizing on sophisticated diffusion models, DiffusionTrend harnesses latents brimming with prior information to encapsulate the nuances of garment details. Throughout the diffusion denoising process, these details are effortlessly merged into the model image generation, expertly directed by a precise garment mask generated by a lightweight and compact CNN. Differing from other approaches, DiffusionTrend sidesteps the necessity for labor-intensive training of diffusion models on extensive datasets. It also dispenses the need for various types of user-unfriendly model inputs. Our experiments demonstrate that, despite lower metric performance, DiffusionTrend delivers a visually convincing virtual try-on experience, all while maintaining the quality and richness of fashion presentation.

### v2

@@ III-C Virtual Try-On Image Reconstruction
TryOnDiffusion [3] - - 13.447 6.964 - - - -
DCI-VTON [4] 0.0530 0.8920 9.130 0.870 0.0443 - 11.800 -
LaDI-VTON [5] 0.0910 0.8760 9.410 0.160 0.0640 0.9060 6.480 0.220
WarpDiffusion [6] 0.0880 0.9850 8.610 - 0.0890 0.9010 9.187 -
OOTDiffusion [7] 0.0710 0.8780 8.810 0.820 0.0450 0.9270 4.200 0.370
StableVITON [8] 0.0732 0.8880 8.233 0.490 0.0388 0.9370 9.940 0.120
IDM-VTON [9] 0.1020 0.8700 6.290 - 0.0620 0.9200 8.640 -
Wear-Any-Way [10] 0.0780 0.8770 8.155 0.780 0.0409 0.9340 11.720 0.330
DiffusionTrend (Ours) 0.0918 0.8592 10.433 0.540 0.0720 0.9172 9.704 0.431
III-D Adaptive Sampling Termination
While DiffusionTrend achieves training-free virtual try-on by injecting garment and background latents at predefined timesteps, we observe that the quality of the generated results is sensitive to the total number of sampling steps $T$, especially in terms of color consistency and structural coherence. To mitigate this issue and enhance robustness across diverse generation conditions, we introduce a sampling termination strategy based on the perceptual color difference metric, CIEDE2000 [58] ($\Delta E_{00}$). This strategy allows the model to dynamically determine the optimal termination step during sampling.
Specifically, we compute the average $\Delta E_{00}$ between the generated image and the target garment image in the garment-masked region after each sampling step. Once the color deviation increases relative to the previous step, it indicates that the generative result has started to deviate from the target appearance. At this point, we halt the sampling process and roll back to the previous intermediate result with the lowest $\Delta E_{00}$ value. This approach requires no additional training or parameter tuning and effectively balances quality and efficiency.
Figure 6: Qualitative comparisons on VITON-HD [11](1st $\sim$ 2nd row) and DressCode [35] (3rd $\sim$ 8th row).
Figure 7: Qualitative results on SHHQ-1.0 [59] model images with garments from DressCode [35] and VITON-HD [11].
IV Experimentation
IV-A Experimental Setup
Dataset. We conduct extensive experiments on two high-resolution datasets from the VITON benchmark: VITON-HD [11] and DressCode [35]. The VITON-HD dataset includes 13,679 pairs of images, each featuring a front-view upper-body shot of women alongside corresponding in-store garments, split into 11,647 training pairs and 2,032 testing pairs. The DressCode dataset is larger, with 48,392 training pairs and 5,400 testing pairs, featuring front-view full-body images of individuals with corresponding in-store garments, categorized into upper-body, lower-body, and dresses.
@@ IV-A Experimental Setup
Evaluation Metrics. We assess both paired and unpaired scenarios. In the paired scenario, the target human image and its corresponding garment image are used for reconstruction. In the unpaired scenario, different garment images are used for the virtual try-on experience. To evaluate the quality of images generated in the paired scenario, we use LPIPS [60] and SSIM [61] metrics to measure resemblance to the original image. In the unpaired scenario, we use FID [62] and KID [63] to gauge the realism and fidelity of the synthesized images.
Implementation Details. We use the Adam optimizer [64] for the apparel localization network with a learning rate of 1e-4, halving it every 10 epochs. The network is trained for 35 epochs on two RTX 3090 GPUs. During inference, the number of clusters $K$ is set to 5. We set the apparel-related prompt to “clothes” and apply our method to Stable Diffusion XL (SDXL) [65]. For the inversion phase, we use an empty prompt, and for model image generation, the prompt is “model wearing clothes.” We conduct DDIM sampling [40] with 50 steps and set the classifier-free guidance to 7.5. Garment latent infusion occurs at time step $t_{1}=40$, and background restoration at $t_{2}=15$. The entire generation process is carried out on a single RTX 3090 GPU.
IV-B Experimental Results
Quantitative Results. Table II shows the quantitative comparisons between DiffusionTrend and other methods on VITON-HD [11] and DressCode [35] test datasets. The results on DressCode of DCI-VTON [4] and StableVITON [8] contain our implementation because no usable codes are given. Despite relying solely on pre-trained models for tasks like garment latent infusion and background restoration, our performance lags behind SOTA models. Past methods incur high training costs to ensure generated try-on images remain realistic and natural under various complex poses, as confirmed by evaluation metrics. In contrast, DiffusionTrend aims to provide a resource-efficient, user-friendly tool for quickly confirming purchase intentions under simple try-on poses. Without training the diffusion model on large datasets, our basic operations fall short in handling complex poses, leading to suboptimal performance. As the first visual try-on model without training on diffusion models, our approach differs from traditional methods in motivation and technical approaches, making traditional try-on dataset evaluations insufficient for comprehensively measuring our method’s effectiveness. Thus, quantitative experiments serve only as a reference, while qualitative experiments will further demonstrate our superiority.
@@ IV-B Experimental Results
Qualitative Results. Fig. 6 provides the qualitative comparison of DiffusionTrend with the state-of-the-art baselines on the VITON-HD [11] and DressCode [35] datasets. The results indicate that our DiffusionTrend performs as well as baseline methods under simple poses.
@@
To further evaluate the generalizability of our method across datasets, we use person images from SHHQ-1.0 [59] as the model image and garments from DressCode [35] and VITON-HD [11] for try-on. As shown in Fig. 7, our method produces visually coherent and naturally blended results, demonstrating stable performance and generalizability across datasets.
@@
TABLE III: Quantitative metrics on virtual try-on datasets for validating mask prediction accuracy.
@@
Our results, as presented in Table III, demonstrate that our compact CNN achieves high accuracy in mask extraction tasks, with IoU scores reaching 0.7 and Dice scores exceeding 0.8. While its accuracy may lag behind SOTA segmentation models, it is more than sufficient for handling virtual try-on tasks. These results validate the effectiveness and practicality of our approach.
IV-C Ablation Study
In this section, we examine the optimality of our method’s components and experimental settings, including background restoration and the timestep for latent infusion or replacement (Sec. III-C), garment mask refinement (Sec. III-B), garment warping and adaptive sampling termination (Sec. III-C, Sec. III-D).
TABLE IV: Quantitative ablations for $t_{1}$ in Garment Latent Infusion and $t_{2}$ in Background Restoration. Bold numbers indicate the best performance.
@@ IV-C Ablation Study
      50       15       0.0791       0.9142       9.8943       0.43
      45       0.0756       0.9153       9.7351       0.43
      40 (Ours)       0.0720       0.9172       9.7040       0.43
      35       0.0712       0.9177       9.8118       0.45
      30       0.0698       0.9183       10.0197       0.47
      25       0.0691       0.9187       10.1865       0.47
      40       25       0.0754       0.9144       9.2217       0.40
      20       0.0737       0.9160       9.4248       0.42
      15 (Ours)       0.0720       0.9172       9.7040       0.43
      10       0.0724       0.9173       10.0531       0.46
      5       0.0727       0.9173       10.6983       0.50
      0       0.1676       0.8527       15.9559       0.73
@@
$t_{1}$ for Garment Latent Infusion. The upper part of Table IV shows that incorporating clothing information too early tends to lower the LPIPS, SSIM, and FID scores. Specifically, as observed in the top row of Fig. 8, prematurely incorporating garment information results in a deficiency of integration between the model and the garment, effectively reconstructing their respective latent representations in disparate areas without any interaction. This issue is clearly visible, as there are pronounced boundaries between the clothing and the human figure, creating an impression of disjunction rather than a cohesive unity. On the other hand, introducing the information at a later stage impedes the development and enhancement of the garment’s finer details. This is observable in the figure, where the cartoon patterns are prone to becoming indistinct, thereby compromising the overall quality of the image.
@@
$t_{2}$ for Background Restoration. Although the quantitative results in Table IV show that the FID and KID scores are higher at $t_{2}=25$ and $20$, it can be observed from the second row of Fig. 8 that, performing background restoration too early can negatively impact the generation of garment details. For instance, at $t_{2}=25$, the waist’s metallic embellishments are lost during the reconstruction process; at $t_{2}=20$, the situation improves slightly, but the metallic dots on either side are still missing. In contrast, performing it too late creates a distinct boundary between the background and the foreground, resulting in unnatural outcomes, and leads to the over-rendering of details. In the figure at $t_{2}=15$, there are small red artifacts visible around the metallic dots.
Garment Mask Refinement Strategy. Initially, to determine an effective approach for separating upper and lower garments from the coarse masks, we evaluate several structural decomposition methods. Region growing tends to over-expand into background areas due to intensity fluctuations, while graph cut often includes the entire human silhouette based on global contrast, deviating from our goal of localized garment segmentation. We observe that a basic K-means clustering can achieve a reasonably effective separation between upper and lower garments. To further improve the segmentation precision and boundary quality, we explore two post-processing techniques: guided filtering and active contour. The former smooths edges but lacks structural modeling ability, while the latter explicitly optimizes contour continuity through energy minimization. Considering both performance and efficiency, we adopt the combination of K-means clustering and active contour refinement as our final strategy. This setup achieves more natural garment boundaries with minimal additional overhead. Qualitative comparisons are presented in Fig. 9.
Figure 9: Visual comparison of different garment structure decomposition strategies.
Garment Warping Strategy. To assess the feasibility of garment warping strategies, we conduct a comparative study on several representative explicit warping modules and perspective transformations. Specifically, we extract the warping modules from: (1) DCI-VTON [4], which represents a dense correspondence-based appearance flow approach; (2) SCW-VTON [55], which introduces shape constraints into the appearance flow to improve contour alignment; (3) PL-VTON [56], which combines affine transformation with dense flow. These approaches represent typical warping modules in diffusion-based and CNN-based virtual try-on pipelines. All of them require dedicated training and rely on structured annotations such as parsing maps and keypoints. To ensure a fair comparison, we integrate their warped garment outputs into the DiffusionTrend pipeline and compared the try-on results with those using perspective transformation. The visual comparison is presented in Fig. 10.
Both visual results and technical analysis reveal several shared drawbacks among these warping methods: (1) Unstable results with distorted edges: Outputs vary considerably across samples, often showing irregular deformations or missing garment parts after warping. (2) Poor generalization and weak robustness: These modules, trained on specific datasets (e.g., VITON [66] or VITON-HD [11]), but exhibit clear structural failures on the DressCode [35] dataset, as shown in our visual examples. (3) Heavy reliance on structural inputs: Most methods require annotations such as keypoints, parsing maps, or densepose, which increases system complexity and reduces accessibility. (4) Strong coupling with downstream try-on models: Although the warping modules are trained independently, their outputs are typically used together with try-on networks during training, leading to joint optimization where the try-on network compensates for warping errors to some extent. However, since we do not train the diffusion model, these imperfect warped results cannot be rectified by the diffusion process itself, leading to failure cases.
Beyond trained warping modules, we also explore non-learning heuristics such as region-based pasting using parsing masks or relative keypoint-based garment stretching. While theoretically simple and training-free, these approaches also rely on structured annotations and showed inferior visual consistency.
In light of these conclusions, we retain perspective transformation as our final garment alignment strategy. Despite its simplicity, it offers consistent and controllable warping results without requiring structural supervision or additional training, making it well-suited to the training-free paradigm of DiffusionTrend.
Figure 10: Comparison of results using different explicit warping modules and perspective transformations. Left: garment with warped/perspective transformation; Right: results generated using the processed garment.
Figure 11: Effectiveness of our adaptive sampling termination across different inversion methods.
Adaptive Sampling Termination. We conduct an experiment in which the number of sampling steps is gradually reduced from 50 to 44, in order to evaluate how this parameter affects the generation quality of DiffusionTrend and to verify the effectiveness of our adaptive sampling termination strategy. As shown in Fig. 11, varying the sampling steps leads to three major types of degradation: (1) Color deviation, where higher sampling steps (e.g., 50) produce oversaturated garments and backgrounds with noticeable reddish or overly bright tones; (2) Loss of detail, where steps below a threshold (e.g., $<45$) result in blurry textures and contours; and (3) Structural distortion, where overly few steps lead to background blending failures and facial deformations.
We attribute this sensitivity to two key factors. First, the training-free nature of DiffusionTrend leaves sampling steps as the only adjustable variable to balance quality and efficiency. Second, a larger number of sampling steps tends to amplify prompt influence, which can cause unwanted deviation from the appearance of the reference garment.
Our adaptive sampling termination strategy mitigates this by terminating sampling at the point where visual fidelity begins to decline. To ensure early-stage stability, this termination strategy is only activated after step $t=45$ in our implementation. As shown in Fig. 11, under the standard SDXL [65] + DDIM Inversion [40] pipeline, our adaptive sampling termination strategy consistently terminates at step 46, aligning well with our empirical observation. We further applied the same strategy to the SDXL + EasyInv [67] setup, where the model automatically selected step 47 as the optimal stopping point. The generated results are visually consistent and stable, demonstrating the robustness and reliability of our strategy across different settings.
V Limitations and Future Work
@@
Reliance on Inversion Quality. Our proposed DiffusionTrend’s performance is indeed influenced by the quality of DDIM inversion results. However, we find that the evolution of more powerful diffusion models is likely to yield better results even with the same inversion method, As demonstrated in Fig. 12. Moreover, ongoing advancements in inversion methods, such as NULL-Text Inversion [68], ReNoise [69], Fixed-Point Iteration [70], and EasyInv [67], provide promising directions for further improvements.
Furthermore, our framework is inherently compatible with ongoing inversion advances and can directly benefit from them without any architectural changes. We thus view current inversion limitations not as constraints, but as evolving components of an improving ecosystem. We believe that the emergence of these improved diffusion models and inversion techniques will significantly expand the application space of our training-free try-on paradigm.
@@
Inference Overhead and Optimization Prospects. DiffusionTrend requires additional time during inference due to the DDIM inversion process, which involves a similar number of steps as DDIM sampling. This time overhead mainly occurs during the inference process and is a potential drawback. However, it can be mitigated through quantization (e.g., Q-Diffusion [71], NDTC [72]) and pruning (e.g., LD-Pruner [73]). While this paper focuses on reducing training overhead, combining these techniques to optimize inference speed will be an important direction for future research.
@@
VI Conclusion
In this paper, we have introduced DiffusionTrend, a novel try-on methodology that forgoes the need for training diffusion models, thereby offering straightforward, conventional pose virtual try-on services with minimal computational demands. Capitalizing on sophisticated diffusion models, DiffusionTrend harnesses latents brimming with prior information to encapsulate the nuances of garment details. Throughout the diffusion denoising process, these details are effortlessly merged into the model image generation, expertly directed by a precise garment mask generated by a lightweight and compact CNN. Differing from other approaches, DiffusionTrend sidesteps the necessity for labor-intensive training of diffusion models on extensive datasets. It also dispenses with the need for various types of user-unfriendly model inputs. Our experiments demonstrate that, despite lower metric performance, DiffusionTrend delivers a visually convincing virtual try-on experience, all while maintaining the quality and richness of fashion presentation.